# CompMarkGS: Robust Watermarking for Compressed 3D Gaussian Splatting

### v5

@@ Abstract
3D Gaussian Splatting (3DGS) is increasingly adopted in various academic and commercial applications due to its real-time and high-quality rendering capabilities, emphasizing the growing need for copyright protection technologies for 3DGS. However, the large model size of 3DGS requires developing efficient compression techniques. This highlights the necessity of an integrated framework that addresses copyright protection and data compression for 3D content. Nevertheless, existing 3DGS watermarking methods significantly degrade watermark performance under 3DGS compression methods, particularly quantization-based approaches that achieve superior compression performance. To ensure reliable watermark detection under compression, we propose a compression-tolerant anchor-based 3DGS watermarking, which preserves watermark integrity and rendering quality. This is achieved by introducing anchor-based 3DGS watermarking. We embed the watermark into the anchor attributes, particularly the anchor feature, to enhance security and rendering quality. We also propose a quantization distortion layer that injects quantization noise during training, preserving the watermark after quantization-based compression. Moreover, we employ a frequency-aware anchor growing strategy that improves rendering quality and watermark performance by effectively identifying Gaussians in high-frequency regions. Extensive experiments demonstrate that our proposed method preserves the watermark even under compression and maintains high rendering quality.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) kerbl20233d has recently emerged as an impactful method for novel view synthesis, offering real-time and photorealistic rendering capabilities. Since its introduction, 3DGS has undergone rapid advancements in the field of image-based 3D reconstruction charatan2024pixelsplat ; lin2024vastgaussian ; zhang2024gs , attracting significant attention from academia and industry. One drawback of 3DGS is that it is composed of a vast number of 3D Gaussians, requiring substantial storage space. This limitation has motivated the development of 3DGS compression techniques to reduce storage requirements and facilitate efficient transmission of 3DGS models.
In parallel, its growing commercial adoption in areas such as digital twin construction and AR/VR applications has raised critical concerns regarding the copyright protection of trained 3DGS models, which has led to increased research interest in 3DGS watermarking. However, to the best of our knowledge, no prior work has jointly tackled model compression and embedding watermarks for 3DGS models. In this paper, we propose the first 3DGS framework that jointly performs watermark embedding and compression, which can be used for ownership verification even after compression, as shown in Fig.1.
@@
Existing 3DGS watermarking methods chen2024guardsplat ; guo2024splats ; huang2025gaussianmarker ; jang20243dgsw3dgaussiansplatting ; tan2024water embed watermarks by directly modifying Gaussian attributes, which makes the watermark vulnerable to loss when those attributes are altered during compression. In particular, quantization-based compression shifts the distribution of model parameters during the compression process, severely degrading watermark performance. As a result, existing methods fail to provide adequate copyright protection for compressed 3DGS models.
To handle this problem, we propose CompMarkGS, an anchor-based 3DGS watermarking strategy. Our method leverages the anchor-based 3DGS architecture, which dynamically predicts Gaussian attributes from anchor points via multiple implicit MLPs. This implicit representation makes it extremely difficult for an adversary to detect the presence of a watermark by directly analyzing the point cloud or Gaussian attributes. We seamlessly integrate the watermark by embedding a learnable watermark feature into the anchor feature, which is one of the anchor attributes, without altering the original model architecture. Unlike prior 3DGS watermarking methods that consider only image-domain distortions, we apply a quantization distortion layer during training to inject quantization noise into the watermark-embedded anchor attributes. By doing so, the embedded watermark learns to adapt to quantization-based compression. As a result, our model preserves watermark integrity through the compression process while maintaining high rendering quality.
To further improve rendering quality, we introduce a frequency-aware anchor growing strategy. Specifically, we identify high-frequency regions in the rendered image, map them to Gaussian coordinates, and selectively apply anchor growing in those regions. This strategy improves both rendering quality and the bit accuracy of the embedded watermark.
Furthermore, we design a HSV loss to address color artifacts that occur when embedding the watermark. To identify the artifacts, we construct a binary mask on the RGB channels by analyzing the hue (H), saturation (S), and value (V) components of the rendered images. This mask guides the loss computation by focusing regions with noticeable color artifacts, improving rendering quality.
Our extensive experiments on the Blender mildenhall2021nerf , LLFF mildenhall2019local , and Mip-NeRF 360 barron2022mip datasets show that CompMarkGS preserves watermark integrity and rendering quality more effectively both before and after compression. Each component of our framework—anchor-based 3DGS watermarking, the quantization distortion layer, and the frequency-aware anchor growing strategy—independently enhances watermark fidelity while preserving rendering quality, even under compression and various watermarking attacks. Furthermore, our method outperforms state-of-the-art approaches across various watermark message lengths. Our main contributions are as follows:
  • We propose CompMarkGS, a compression-tolerant anchor-based watermarking framework for 3DGS that embeds a learnable watermark embedding feature into anchor features, preserving the model structure and ensuring high rendering quality and bit accuracy.
@@
  • We propose a frequency-aware anchor growing strategy that selectively expands anchors in high-frequency regions, which enhances both rendering quality and watermark bit accuracy, and a HSV loss that mitigates color artifacts to further improve rendering quality.
@@ 2 Related Works
3D Gaussian Splatting (3DGS) kerbl20233d is an innovative technique for 3D representation that employs explicit primitives to model 3D worlds. It renders images by projecting these primitives into arbitrary viewpoints and seamlessly blending pixel colors via an $\alpha$-blender, achieving high-quality results while substantially accelerating real-time rendering. Thanks to these advantages, 3DGS has been widely adopted across various research areas, including avatars abdal2024gaussian ; chen2025taoavatar ; qian20243dgs ; yuan2024gavatar , dynamic scenes li2024spacetime ; wu20244d ; yan2025instant ; zhu2025motiongs , and 3D generation tang2023dreamgaussian ; xie2024physgaussian ; yi2024gaussiandreamer ; zhou2024diffgs . Recently, Scaffold-GS lu2024scaffold has advanced the 3DGS framework by introducing anchor points to construct a hierarchical 3D representation. Anchor-based methods lee2024mode ; ren2024octree ; wang2024learning effectively minimize redundant Gaussians, enhancing rendering quality and increasing robustness against view changes. This representation approach is promising for 3DGS compression optimization, as it reduces parameter counts while maintaining high-quality rendering.
@@
To achieve high-quality rendering images, 3DGS kerbl20233d generates a considerable number of Gaussians, leading to significant storage overhead. To address the storage overhead, vector quantization methods fan2025lightgaussian ; lee2024compact ; navaneet2024compgs ; xie2024mesongs have been widely explored. These methods prune Gaussians with minimal impact on rendering quality and use codebooks to encode the attributes of Gaussians compactly. More recently, anchor-based representations have gained significant attention in 3DGS chen2024hac ; lu2024scaffold ; wang2024contextgs . Compression methods chen2025pcgs ; chen2024hac ; wang2024contextgs ; zhan2025cat that leverage the structural relationships between anchors have shown superior performance. For example, HAC chen2024hac reduces spatial redundancies among anchors by utilizing hash grids for parameter quantization and enabling entropy modeling. ContextGS wang2024contextgs introduces a unified compression framework with a factorized prior, enabling entropy modeling of anchor features while leveraging hierarchical anchor relations to reduce redundancy among anchors. CAT-3DGS zhan2025cat projects Gaussian primitives onto PCA-aligned triplanes and applies spatial autoregressive coding to capture spatial correlations, thereby enhancing entropy coding efficiency. Given their compression efficiency, anchor-based representations are especially advantageous for real-world applications.
Digital watermarking.
Digital watermarking has long been studied to protect the copyright of digital assets. Early approaches embed information in the pixel domain van1994digital ; wolfgang1996watermark or the frequency domain barni2001improved ; kutter1997digital ; navas2008dwt ; urvoy2014perceptual . Driven by advances in deep learning, encoder–decoder–based methods jiang2023evading ; luo2020distortion ; sander2024watermark ; zhu2018hidden were proposed, demonstrating competitive robustness against a variety of distortions compared to traditional techniques. Recently, efforts have extended watermarking to protect 3D models and their rendered content zhang2023adaptive ; article . In particular, in radiance field model watermarking, methods have been proposed to embed watermarks by leveraging the implicit neural representation (INR) of NeRF mildenhall2021nerf . CopyRNeRF luo2023copyrnerf is the first digital watermarking method for implicit NeRF, introducing a watermarked color representation. WateRF jang2024waterf is a plug-and-play method compatible with all radiance field representations, leveraging the frequency domain to achieve robust watermarking. With the advent of 3DGS kerbl20233d , various watermarking methods have been proposed. GS-Hider zhang2025gs was the first to introduce steganography for 3DGS, and SecureGS zhang2025securegs proposes an anchor-based steganographic framework. GaussianMarker huang2025gaussianmarker embeds the watermark by estimating the uncertainties of Gaussians and incorporating watermark-specific Gaussians into the existing model based on these uncertainties. 3D-GSW jang20243dgsw3dgaussiansplatting achieves watermark embedding by selectively pruning Gaussians with minimal impact on rendering quality. GuardSplat chen2024guardsplat introduces an SH-aware message-embedding mechanism. However, model compression often leads to watermark loss, and robust watermarking methods against compression remain underexplored. Therefore, our work integrates 3DGS training with watermark embedding, ensuring robustness against degradations introduced by model compression.
@@ 3 Preliminaries
Scaffold-GS lu2024scaffold clusters adjacent Gaussians using anchor points, reducing redundancy. Initial anchor points are placed at voxel centers, and the attributes for each anchor consist of an anchor feature $f\in\mathbb{R}^{d}$, scaling factor $l\in\mathbb{R}^{3}$, and $K$ learnable offsets ${O}\in\mathbb{R}^{K\times 3}$, where $K$ is the number of Gaussians. Visible anchor points within the viewing frustum generate $K$ Gaussians, with their positions computed as follows:
@@
Figure 2: Overview of CompMarkGS. We obtain the watermarked anchor feature $f^{w}$ by adding the watermark embedding feature $f^{\prime}$ to the anchor feature $f$. Quantized watermarked anchor feature from the quantization distortion layer are fed into individual MLPs to predict Gaussian attributes. We input the rendered image’s low-frequency band extracted via DWT into a pre-trained decoder to extract the watermark. Frequency-aware anchor growing strategy identifies Gaussians in high-frequency regions to introduce additional anchors, while an HSV loss $\mathcal{L}_{hsv}$ and a frequency loss $\mathcal{L}_{freq}$ further enhance rendering quality. Finally, the entire model is optimized using a total loss $\mathcal{L}_{total}$.
@@ 4.1 Anchor-based 3DGS watermarking
Anchor-based 3DGS lu2024scaffold , enhances the security of embedded watermarks by hiding explicit Gaussian attributes behind anchor features processed by implicit MLPs. This property of anchor-based 3DGS makes it more difficult for attackers to directly manipulate embedded watermarks and Gaussian attributes. Moreover, the anchor feature $f$ influences the rendering quality indirectly through the MLPs rather than directly controlling Gaussian attributes. On the other hand, attributes like the position, scaling factor $l$, and offsets $O$ directly determine Gaussian attributes, and modifying them leads to visible distortions in the rendered images.
To ensure rendering quality, we embed the watermark into the anchor feature, introducing a learnable watermark embedding feature $f^{\prime}\in\mathbb{R}^{d}$ of the same dimensionality as the anchor feature $f\in\mathbb{R}^{d}$. The resulting watermarked anchor feature $f^{w}$ is defined as follows:
@@
where $\mathrm{tanh}(\cdot)$ represents the element-wise hyperbolic tangent function. To prevent the watermark embedding feature from growing excessively, we use $\mathrm{tanh}(\cdot)$ to constrain its values to the range of $[-1,1]$ before adding it to the anchor feature.
@@ 4.2 Robustness to quantization-based compression
We aim to ensure that the watermarked model retains its watermark information after the model compression. Since many compression methods rely on quantization, we incorporate a quantization distortion layer to improve watermark robustness during compression. Specifically, we leverage the quantization technique used in anchor-based 3DGS compression chen2024hac to inject quantization noise into the watermarked anchor feature $f^{w}$ during training. This approach prevents the loss of watermark information during subsequent quantization-based compression. The quantized watermarked anchor feature $\tilde{f}^{w}_{i}$ for the $i$-th anchor is defined as follows:
@@
where $q_{i}\in\mathbb{R}$ denotes the quantization scale restricted to $[0,2\mathrm{Q_{0}}]$, and $r_{i}\in\mathbb{R}$ is the refinement output from an $\text{MLP}_{q}$ that takes $f^{w}_{i}$ as input, used to adjust the initial quantization scale $\mathrm{Q_{0}}$. Thus, by scaling a randomly generated $d$-dimensional noise vector from a uniform distribution over $[-\frac{1}{2},\frac{1}{2}]$ by $q_{i}$ and adding it to the watermarked anchor feature $f^{w}$, we simulate the rounding error caused by quantization, thus reflecting the effects of quantization during training.

### v6

@@ Abstract
As 3D Gaussian Splatting (3DGS) is increasingly adopted in various academic and commercial applications due to its high-quality and real-time rendering capabilities, the need for copyright protection is growing. At the same time, its large model size requires efficient compression for storage and transmission. However, compression techniques, especially quantization-based methods, degrade the integrity of existing 3DGS watermarking methods, thus creating the need for a novel methodology that is robust against compression. To ensure reliable watermark detection under compression, we propose a compression-tolerant 3DGS watermarking method that preserves watermark integrity and rendering quality. Our approach utilizes an anchor-based 3DGS, embedding the watermark into anchor attributes, particularly the anchor feature, to enhance security and rendering quality. We also propose a quantization distortion layer that injects quantization noise during training, preserving the watermark after quantization-based compression. Moreover, we employ a frequency-aware anchor growing strategy that enhances rendering quality by effectively identifying Gaussians in high-frequency regions, and an HSV loss to mitigate color artifacts for further rendering quality improvement. Extensive experiments demonstrate that our proposed method preserves the watermark even under compression and maintains high rendering quality.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has recently emerged as an impactful method for novel view synthesis, offering real-time and photorealistic rendering capabilities. Since its introduction, 3DGS has undergone rapid advancements in the field of image-based 3D reconstruction (Charatan et al., 2024; Lin et al., 2024; Zhang et al., 2024), attracting significant attention from academia and industry. One drawback of 3DGS is that it is composed of a vast number of 3D Gaussians, requiring substantial storage space. This limitation has motivated the development of 3DGS compression techniques to reduce storage requirements and facilitate efficient transmission of 3DGS models. In parallel, its growing commercial adoption in areas such as digital twin construction and AR/VR applications has raised critical concerns regarding the copyright protection of trained 3DGS models. These concerns have led to increased research interest in 3DGS watermarking. However, to the best of our knowledge, no prior work has specifically addressed the challenge of creating a 3DGS watermark that can withstand model compression, a scenario illustrated in Fig. 1.
Existing 3DGS watermarking methods (Chen et al., 2024b; Guo et al., 2024; Huang et al., 2025; Jang et al., 2025; Tan et al., 2024) embed watermarks by directly modifying Gaussian attributes, which makes the watermark vulnerable to loss when those attributes are altered during compression. In particular, quantization‐based compression shifts the distribution of model parameters during the compression process, severely degrading watermark performance. As a result, existing methods fail to provide adequate copyright protection for compressed 3DGS models.
@@
To handle this problem, we propose CompMarkGS, an anchor-based 3DGS watermarking method. Our method leverages the anchor-based 3DGS architecture, which dynamically predicts Gaussian attributes from anchor points via multiple implicit multi-layer perceptrons (MLPs). This implicit representation makes it extremely difficult for a malicious user to detect the presence of a watermark by directly analyzing the point cloud or Gaussian attributes. We seamlessly integrate the watermark by embedding a learnable watermark feature into the anchor feature, which is one of the anchor attributes, without altering the original model architecture. Unlike prior work focusing on image-domain distortions, we tackle the novel challenge of watermark robustness against model quantization. We propose a quantization distortion layer (QDL) that injects quantization noise into anchor attributes during training, teaching the watermark to be resilient. As a result, our model achieves both high rendering quality and robust watermark integrity after compression.
To further improve rendering quality, we introduce a frequency-aware anchor growing strategy. Specifically, we identify high-frequency regions in the rendered image, map them to Gaussian coordinates, and selectively apply anchor growing in those regions. In addition to this structural enhancement, we design an HSV loss to address color artifacts that occur when embedding the watermark. This loss function operates by constructing a binary mask over regions with noticeable color artifacts, identified by analyzing the hue (H), saturation (S), and value (V) components. This mask guides the loss computation by focusing on these specific regions, improving the rendering quality.
Our extensive experiments on the Blender (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022) datasets show that CompMarkGS preserves watermark integrity and rendering quality more effectively both before and after compression. Each component of our method—anchor-based 3DGS watermarking, the quantization distortion layer, and the frequency-aware anchor growing strategy—independently enhances watermark fidelity while preserving rendering quality, even under compression and various watermarking attacks. Furthermore, our method outperforms state-of-the-art approaches across various watermark message lengths. Our main contributions are as follows:
  • We propose CompMarkGS, a compression-tolerant anchor-based watermarking method for 3DGS that embeds a learnable watermark embedding feature into the anchor feature, preserving the model structure while ensuring high rendering quality and bit accuracy.
@@
  • We propose a frequency-aware anchor growing strategy that selectively expands anchors in high-frequency regions, which enhances rendering quality, and an HSV loss that mitigates color artifacts to further improve rendering quality.
@@ 2 Related Works
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) is an innovative 3D representation technique that models scenes using explicit primitives. It renders images by projecting these primitives into arbitrary viewpoints and seamlessly blending pixel colors via an $\alpha$-blending, achieving high-quality results while substantially accelerating real-time rendering. Thanks to these advantages, 3DGS has been widely adopted in various research areas, including avatars (Abdal et al., 2024; Chen et al., 2025a; Qian et al., 2024; Yuan et al., 2024), dynamic scenes (Li et al., 2024; Wu et al., 2024; Yan et al., 2025; Zhu et al., 2025), and 3D generation (Tang et al., 2023; Xie et al., 2024b; Yi et al., 2024; Zhou et al., 2024). Recently, Scaffold-GS (Lu et al., 2024) has advanced the 3DGS framework by introducing anchor points to construct a hierarchical 3D representation. Anchor-based methods (Lee et al., 2024b; Ren et al., 2024; Wang et al., 2024b) effectively minimize redundant Gaussians, enhancing rendering quality and increasing robustness against view changes. This representation approach is promising for 3DGS compression optimization, as it reduces parameter counts while maintaining high-quality rendering.
@@
To achieve high-quality rendering images, 3DGS generates a considerable number of Gaussians, leading to significant storage overhead. To address the storage overhead, vector quantization methods (Fan et al., 2025; Lee et al., 2024a; Navaneet et al., 2024; Xie et al., 2024a) have been widely explored. These methods prune Gaussians with minimal impact on rendering quality and use codebooks to encode the attributes of Gaussians compactly. More recently, anchor-based representations have gained significant attention in 3DGS (Chen et al., 2024a; Lu et al., 2024; Wang et al., 2024a). Compression methods (Chen et al., 2025b; 2024a; Wang et al., 2024a; Zhan et al., 2025) that leverage the structural relationships between anchors have shown superior performance. For example, HAC (Chen et al., 2024a) reduces spatial redundancies among anchors by utilizing hash grids for parameter quantization and enabling entropy modeling. ContextGS (Wang et al., 2024a) introduces a unified compression framework with a factorized prior, enabling entropy modeling of anchor features while leveraging hierarchical anchor relations to reduce redundancy among anchors. CAT-3DGS (Zhan et al., 2025) projects Gaussian primitives onto PCA-aligned triplanes and applies spatial autoregressive coding to capture spatial correlations, thereby enhancing entropy coding efficiency. Given their compression efficiency, anchor-based representations are especially advantageous for real-world applications.
Steganography and digital watermarking.
Information hiding research is primarily divided into two categories: digital watermarking and steganography, which differ in their core objectives. Steganography aims to conceal information in digital assets with invisibility as the primary metric. Driven by advances in deep learning, various methods (Tancik et al., 2020; Biswal et al., 2024) have been proposed to hide information in media such as images and videos, with recent work (Dong et al., 2024; Li et al., 2023) extending into 3D scene representation. For example, GS-Hider (Zhang et al., 2025a) introduces the first steganography technique for 3D Gaussian Splatting (3DGS), and SecureGS (Zhang et al., 2025b) proposes an anchor-based steganographic framework that hides information using a private MLP. In contrast, digital watermarking embeds ownership data to protect assets and prioritizes robustness to distortions. Early watermarking research focused on the pixel domain (Van Schyndel et al., 1994; Wolfgang & Delp, 1996) or the frequency domain (Barni et al., 2001; Navas et al., 2008). Subsequently, deep learning-based methodologies (Luo et al., 2020; Zhu et al., 2018) demonstrated superior robustness against a variety of distortions. This research has recently expanded to protect radiance field models like NeRF (Mildenhall et al., 2021). WateRF (Jang et al., 2024) proposes a plug-and-play method that uses the frequency domain for robust watermarking. With the advent of 3DGS, GaussianMarker (Huang et al., 2025) embeds the watermark by estimating Gaussian uncertainties and integrating specific Gaussians. 3D-GSW (Jang et al., 2025) embeds the watermark by selectively removing Gaussians to minimize the impact on rendering quality. GuardSplat (Chen et al., 2024b) introduces an SH‐aware embedding mechanism. However, these works only consider traditional distortions and do not account for real-world attacks such as model compression. In this paper, we propose a watermarking method that is robust to model compression.
@@ 3 Preliminaries
Scaffold-GS (Lu et al., 2024) clusters adjacent Gaussians using anchor points, reducing redundancy. Initial anchor points are placed at voxel centers, and the attributes for each anchor consist of an anchor feature $f\in\mathbb{R}^{d}$, scaling factor $l\in\mathbb{R}^{3}$, and $K$ learnable offsets ${O}\in\mathbb{R}^{K\times 3}$, where $K$ is the number of Gaussians. Visible anchor points within the viewing frustum generate $K$ Gaussians, with their positions computed as follows:
@@
Figure 2: Overview of CompMarkGS. A learnable watermark embedding feature $f^{w}$f is added to the anchor feature $f$ to create a watermarked anchor feature $f^{w}$. The quantized watermarked anchor feature from the quantization distortion layer is then fed into individual MLPs to predict Gaussian attributes. The watermark is extracted from the low-frequency band of the rendered image via a pre-trained decoder. A frequency-aware anchor growing strategy identifies Gaussians in high-frequency regions to introduce additional anchors. Moreover, an HSV loss $\mathcal{L}_{hsv}$, and a frequency loss $\mathcal{L}_{freq}$, are applied to mitigate visual artifacts. The entire model is optimized via the total loss $\mathcal{L}_{total}$.
@@ 4.1 Anchor-based 3DGS watermarking
Anchor-based 3DGS (Lu et al., 2024) enhances the security of embedded watermarks by hiding explicit Gaussian attributes behind anchor features processed by implicit MLPs. This property of anchor-based 3DGS makes it more difficult for attackers to directly manipulate embedded watermarks and Gaussian attributes. Moreover, the anchor feature $f$ influences the rendering quality indirectly through MLPs rather than directly controlling Gaussian attributes. On the other hand, attributes such as position $x$, scaling factor $l$, and offsets $O$ directly determine Gaussian attributes, and modifying them leads to visible distortions in the rendered images. To embed a watermark into the anchor feature while preserving rendering quality, we introduce a learnable watermark embedding feature $f^{\prime}\in\mathbb{R}^{d}$, with the same dimension as the anchor feature $f\in\mathbb{R}^{d}$. This design aims to achieve an optimal balance between rendering quality and watermark robustness. The resulting watermarked anchor feature $f^{w}$ is defined as follows:
@@
where the watermarked anchor feature $f^{w}$ is generated by the element-wise sum of the anchor feature $f$ and a learnable watermark embedding feature $f^{\prime}$. The term $\mathrm{tanh}(\cdot)$ denotes the element-wise hyperbolic tangent function, which we apply to the watermark embedding feature $f^{\prime}$ to bound its values within the range [1,1][-1,1] before addition. This bounding process is crucial for stabilizing gradients and preserving rendering quality. In a well-trained anchor-based 3DGS model, the components of the anchor attributes are known to follow a normal distribution (Lu et al., 2024). Adding an unbounded feature would increase the variance of this distribution, leading to unstable gradients, slower convergence, and degraded rendering quality. While a sigmoid function could be used, its $[0,1]$ output range would introduce a directional bias by only adding positive values. The symmetric [1,1][-1,1] range of the tanh function prevents this bias, ensuring the watermark is added with a consistent magnitude. As a result, our method achieves stable convergence and efficient embedding without degrading rendering quality. A detailed analysis is provided in Appendix E.1.
@@ 4.2 Robustness to quantization-based compression
In digital watermarking research, differentiable distortion layers have long been used to enhance robustness by simulating traditional distortions, such as JPEG and cropping, during training (Zhu et al., 2018). However, this method has not been applied in existing 3DGS watermarking research to the problem of model compression, leading to significant watermark loss during quantization.
To overcome this limitation, we propose a quantization distortion layer (QDL) that simulates quantization-induced distortions, encompassing losses in watermark robustness and rendering fidelity. Our QDL leverages quantization techniques from prior anchor-based 3DGS compression methods (Chen et al., 2024a). While the original method uses this mechanism for model compression, our work is the first to adapt it as a form of data augmentation to train a robust watermark. During watermark training, the QDL simulates the quantization process by injecting noise into the watermarked anchor feature $f^{w}$. By integrating the QDL into the embedding stage, the watermark is thus optimized to remain robust against subsequent compression attacks. The quantized watermarked anchor feature for the $i$-th anchor $\tilde{f}^{w}_{i}$ is defined as follows:
@@
where $q_{i}\in\mathbb{R}$ denotes the quantization scale restricted to $0,2\mathrm{Q_{0}}]$, and $r_{i}\in\mathbb{R}$ is the refinement output from an $\text{MLP}_{q}$ that takes $f^{w}_{i}$ as input to adjust the initial quantization scale $\mathrm{Q_{0}}$. By scaling a randomly generated $d$-dimensional noise vector from a uniform distribution over $\frac{1}{2},\frac{1}{2}]$ by $q_{i}$ and adding it to the watermarked anchor feature $f^{w}$, we simulate the rounding errors caused by quantization. This process forces the watermark to become resilient to the specific distortions it will encounter during a real compression attack.